Determine by Lagrange’s formula, the percentage number of criminals under 35 years:

Age % number of criminals

under 25 years 52

under 30 years 67.3

under 40 years 84.1

under 50 years 94.4


1
Expert's answer
2021-01-12T01:15:30-0500

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 x & 25 & 30 & 40 & 50\\ \hline
y& 52 & 67.3 & 84.1 & 94.4 \\
 \hdashline
 
\end{array}", where "x" denotes age and "y" denotes % of criminals.


By Lagrange’s interpolation formula we have:

"f(x)= \frac{(x-x_1)(x-x_2)(x-x_3)}{(x_0-x_1)(x_0-x_2)(x_0-x_3)}y_0+ \frac{(x-x_0)(x-x_2)(x-x_3)}{(x_1-x_0)(x_1-x_2)(x_1-x_3)} y_1+ \frac{(x-x_0)(x-x_1)(x-x_3)}{(x_2-x_0)(x_2-x_1)(x_2-x_3)}y_2+ \frac{(x-x_0)(x-x_1)(x-x_2)}{(x_3-x_0)(x_3-x_1)(x_3-x_2)} y_3"


We put "x=35:"

"y(35)=f(35)= \frac{(35-30)(35-40)(35-50)}{(25-30)(25-40)(25-50)}52+ \frac{(35-25)(35-40)(35-50)}{(30-25)(30-40)(30-50)} 67.3+ \frac{(35-25)(35-30)(35-50)}{(40-25)(40-30)(40-50)}84.1+ \frac{(35-25)(35-30)(35-40)}{(50-25)(50-30)(50-40)} 94.4=-\tfrac{1}{5}\cdot 52+\tfrac{3}{4}\cdot 67.3+\tfrac{1}{2}\cdot 84.1-\tfrac{1}{20}\cdot 94.4=77.405"



Answer: the percentage number of criminals under 35 years is 77.405.